OpenStudy (anonymous):

if cos(2x-1)=sin(3x+6) then the value of x is? Explanation please

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

cos(2x-1)=sin(3x+6) cos(2x-1)=cos(pi/2 - (3x+6)) cos(2x-1)=cos(pi/2 - 3x - 6) arccos[ cos(2x-1) ] = arccos[ cos(pi/2 - 3x - 6) ] 2x-1 = pi/2 - 3x - 6 now solve for x

OpenStudy (anonymous):

The pi/2 is confusing me, i can't solve for x with that there

jimthompson5910 (jim_thompson5910):

ok let z = pi/2

jimthompson5910 (jim_thompson5910):

2x-1 = pi/2 - 3x - 6 2x-1 = z - 3x - 6 now solve for x

jimthompson5910 (jim_thompson5910):

when you're done, replace z with pi/2

OpenStudy (anonymous):

is it pi/2-1?

jimthompson5910 (jim_thompson5910):

2x-1 = z - 3x - 6 2x = z - 3x - 6+1 2x+3x = z - 6+1 5x = z - 5 x = z/5 - 5/5 x = z/5 - 1 x = (pi/2)/5 - 1 x = pi/10 - 1 So \[\Large x = \frac{\pi}{10} - 1\]

OpenStudy (anonymous):

That isn't one of the choices though

jimthompson5910 (jim_thompson5910):

what are your choices

OpenStudy (anonymous):

the choices are -7, 7, 17, 35

jimthompson5910 (jim_thompson5910):

does it say anything about rounding to the nearest whole number?

OpenStudy (anonymous):

No, but after the parenthesis there are degree signs but i don't think that has any relevance

jimthompson5910 (jim_thompson5910):

oh degrees, not radians

jimthompson5910 (jim_thompson5910):

x = z/5 - 1 x = 90/5 - 1 x = 18 - 1 x = 17

OpenStudy (anonymous):

oh so it did matter, sorry about that, but thank you, ill try to understand it now!

jimthompson5910 (jim_thompson5910):

you're welcome